what is the answer to 5 radical 32 ?
Found 2 solutions by KMST, stanbon:
Answer by KMST(5328)
Answer by stanbon(75887)

what is the answer to 5 radical 32
-------
5*sqrt(16*2) = 5*4*sqrt(2) = 20sqrt(2)
-------------------
Cheers,
Stan H.
==============